The Eastern Shuttle, Inc., is a regional airline providing shuttle service between New York and
Washington, DC. An analysis of the monthly demand for service has revealed the following
demand relation:
Q = 26,000 – 500P – 250POG + 200IB – 5,000S
Where: Q is quantity measured by the number of passengers per month, P is price ($), POG is
a regional price index for other consumer goods (1967 = 1.00), IB is an index of business activity,
and S, a binary or dummy variable, equals 1 in summer months and 0 otherwise.
A. Determine the demand curve facing the airline during the winter month of January if POG
= 4 and IB = 250.
B. Determine the demand curve facing the airline, quantity demanded, and total revenues
during the summer month of July if P = $100 and all other price-related and business
activity variables are as specified previously.
1
Expert's answer
2017-10-12T02:59:07-0400
A. Demand curve is:
Q = 26,000 – 500P – 250*4 + 200*250 – 5,000*0 = 75,000 – 500P
B. Demand curve is:
Q = 26,000 – 500P – 250*4 + 200*250 – 5,000*1 = 70,000 – 500P
Quantity demanded is:
Q = 70,000 – 500 *100 = 20,000 (passengers per month)
Total revenue is:
TR = P*Q = 100*20,000 = 2,000,000 ($ per month)